OpenStudy (anonymous):

8 in = 0.5 g(/\t)^2 How do I solve this ?

OpenStudy (anonymous):

you want to solve for t right? well get t on its own 8 = 0.5 g t^2 8/(0.5g) = t^2 sqrt(4/g) = t

OpenStudy (mrnood):

If your question means the distance is 8 inches then you need to be careful with units. g = 9.81 m/s^2 g= 32 ft/s^2 g= 386 in/s^2 Make sure you use consistent units

OpenStudy (anonymous):

a good point well made

OpenStudy (anonymous):

Would I plug in 368 in/s^2 for g ?